A solution of a metal ion when treated with KI gives a red precipitate which dissolves in excess KI to give a colourless solution. Moreover, the solution of metal ion on treatment with a solution of cobalt (II) thiocyanate gives rise to a deep blue crystalline precipitate.  The metal ion is
a. Pb^2+
b. Hg^2+
c. Cu^2+
d. Co^2+

Grade:12th pass

1 Answers

bharat bajaj IIT Delhi
9 years ago
Hg (2+) + KI = HgI2 ( Red ppt.)
HgI2 + KI (excess) = K2HgI4

HgI^(2+) + Co(SCN)2 = Hg(SCN)2 Blue crystalline precipitate
B) is the answer
THanks
Bharat Bajaj
askiitians faculty